A charge Q is uniformly distributed throughout a sphere of radius a. Taking the potential at infinity as zero, the potential at r = b < a is


1
Expert's answer
2021-07-05T16:07:57-0400

Gauss law

"\smallint E.A=\frac{Q}{\epsilon_0}"

"r=b<a"

"E.(4\pi b^2)=\frac{\rho\frac{4}{3}\pi b^3}{\epsilon_0}"

Where

"\rho=\frac{Q}{\frac{4}{3}\pi a^3}"

"E_{in}=\frac{qb}{4\pi\epsilon_0a^3}"

b>a

"E_{o}.A=\frac{q}{\epsilon_0},E_{o}=\frac{q}{4\pi \epsilon _0b^3}"

Potential

"V_a-V_{infinite}=\smallint_{inf}^{a}E.dr"

"V_a=-\smallint_{inf}^{a}E.dr"

"V_a=-[\smallint _{inf}^bE_o.dr+\int_b^aE_{in}.dr]"

"V_a=-[\smallint _{inf}^b\frac{kQ}{r^2}.dr+\int_b^a\frac{kQr}{R^3}.dr]"

"V_{in}=\frac{kQ(3a^2-b^2)}{2a^3}"

b=a

"V_{in}=\frac{3Q}{8\pi \epsilon_0a}"

"\u2206V=V_{in}{(r=a)}-V_{in}{(r=0)}"

"∆V=\frac{-Q}{8\pi\epsilon_0 a}"